OpenStudy (mydilemma):

Can someone help me understand this? "the midpoint of uv is (5, -11). The coordinates of one endpoint are U(3, 5) find the coordinates of endpoint v"

OpenStudy (misssunshinexxoxo):

y=mx+b y=ΔyΔxx+b y=y2−y1x2−x1x+b y=(−11)−55−3x+b y=−162x+b y=−8x+b substitute in a point to figure out b (I'll put U in). 5=(−8)(3)+b 5=−24+b b=29 So the equation of you're line is: y=−8x+29 Now the symmetry part: The midpoint is +2 in the x direction, so add another 2 to the midpoint to get V's x coordinate. y3=−8x3+29 y3=−8(7)+29 y3=−56+29 y3=−27 V should be (7, -27) s=(Δy)2+(Δx)2−−−−−−−−−−−−√ s=(y2−y1)2+(x2−x1)2−−−−−−−−−−−−−−−−−−√ s=(−11−5)2+(5−3)2−−−−−−−−−−−−−−−−−√ s=(−16)2+(2)2−−−−−−−−−−−√ s=256+4−−−−−−√ s=260−−−√ midpoint to V s=(Δy)2+(Δx)2−−−−−−−−−−−−√ s=(y3−y2)2+(x3−x2)2−−−−−−−−−−−−−−−−−−√ s=((−27)−(−11))2+(7−5)2−−−−−−−−−−−−−−−−−−−−−−√ s=(−16)2+(2)2−−−−−−−−−−−√ s=256+4−−−−−−√ s=260−−−√

OpenStudy (misssunshinexxoxo):

If you have any questions here to help!

OpenStudy (mydilemma):

Okay thank you so much!(:

OpenStudy (mrnood):

 WOW That seems a complicated explanation for a simple concept The midpoint is middle of UV in both X and Y direction SO Find how far to move from end poit to midpoint in x then add the same amount again Find out how far to move from endpoint ot midpoint in Y then add the same amount again. (Note value to 'add' may be negative)